3. Calculate the present value of the following sequence of willingness to pay amounts: this year, $100; next year, $150; year 2, $150; year 3, $50. Use a 5 percent discount rate. Recalculate using a discount rate of 8 per- cent. What is the effect of using a higher discount rate?
Answer #1

r = 5%
PV = 100 + 150/1.05 + 150/1.052 + 50/1.053 = $ 422.1

r = 8%
PV = 100 + 150/1.08 + 150/1.082 + 50/1.083 = $ 407.2

Present value varies inversely with the discount rate (higher discount rate means lower present value)